OpenStudy (anonymous):

(x+3)2 = -13 can somebody please help me

OpenStudy (anonymous):

i need solution set

OpenStudy (anonymous):

\[^{?}\]

OpenStudy (anonymous):

2(x+3)= -13 2x+6 = -13 2x= -13 -6 2x/ 2 = -19/2 x = 9.5

OpenStudy (anonymous):

x= -9.5

OpenStudy (anonymous):

thank u very muchh

OpenStudy (anonymous):

welcome